OpenStudy (anonymous):

how do u solve y=2x+5 and y=-x+2

OpenStudy (anonymous):

Are those as a simultaneous equation? Or separate?

OpenStudy (amistre64):

"and" has a precise mathematical term .. so the solution set is the point of intersection of the lines

OpenStudy (amistre64):

since y = y, just equate them and solve for x once you know x, either equation will give you y

OpenStudy (anonymous):

So its simultaneous then? Got a bit confused there.

OpenStudy (anonymous):

Ok so since both of the equations have y= in front of them, u can just equal them 2 each other like: 2x+5=-x+2 [and now we solve, add x to both sides to get:] 3x+5=2 [Subtract 5 on both sides 2 get:] 3x=-3 [Divide both sides by 3] x=-1